# Take Care of Your Prompt Bias! Investigating and Mitigating Prompt Bias in Factual Knowledge Extraction

### v1

@@ Abstract
Recent research shows that pre-trained language models (PLMs) suffer from “prompt bias” in factual knowledge extraction, i.e., prompts tend to introduce biases toward specific labels. However, the extent and impact of prompt bias within the model remain underexplored. In response, this paper quantifies the bias with various types of prompts and assesses their impact on different benchmarks. We show that: 1) all prompts in the experiments exhibit non-negligible bias, with gradient-based prompts like AutoPrompt and OptiPrompt displaying significantly higher levels of bias; 2) prompt bias can amplify benchmark accuracy unreasonably by overfitting the test datasets, especially on imbalanced datasets like LAMA. Based on these findings, we propose a representation-based approach to mitigate the prompt bias during inference time. Specifically, we first estimate the biased representation using prompt-only querying, and then remove it from the model’s internal representations to generate the debiased representations, which are used to produce the final debiased outputs. Experiments across various prompts, PLMs, and benchmarks show that our approach can not only correct the overfitted performance caused by prompt bias, but also significantly improve the prompt retrieval capability (up to 10% absolute performance gain). Our findings shed new light on the underlying predicting mechanisms of prompt-based queries in PLMs. Hopefully, our plug-and-play approach can be a golden standard to strengthen PLMs toward reliable knowledge bases. Code and data are released in https://github.com/FelliYang/PromptBias.
@@
[…]
${}^{1}$nstitute of Artificial Intelligence, School of Computer Science, Wuhan University, China
${}^{2}$chool of Mathematics and Statistics, Wuhan University, China
${}^{6}$ubei Key Laboratory of Computational Science, Wuhan University, China
${}^{3}$eihang University, ${}^{4}$he University of Sydney, ${}^{5}$anyang Technological University
@@ 1.   Introduction
Although prompt-based querying achieves decent performance, researchers found that it will introduce prompt bias when extracting factual knowledge Cao et al. (2021), as illustrated in Figure 1. While the previous study has demonstrated that prompt bias undermines the reliability of benchmarks, the extent of bias among different prompts and how much impact it will have on benchmark reliability remains unclear.
To this end, this paper proposes a method to quantify the prompt bias and thoroughly assess its impact on the reliability of widely used benchmarks such as LAMA. Experiments show that all knowledge-extraction prompts have significant bias, with previously reported high-performing prompts such as AutoPrompt and Optiprompt often exhibiting more pronounced bias. Moreover, we find that prompt bias severely compromises benchmark reliability by overfitting the test datasets and amplifying performance unreasonably. For example, the prompt bias of OptiPrompt helps inflate absolute accuracy by over 16.47% on the LAMA benchmark when probing the BERT-base model. Additionally, prompt bias can mislead language models and discourage models from making correct predictions, as illustrated in Figure 1. This substantially hinders the knowledge retrieval capabilities of prompts.
To deal with the risk of prompt bias, we propose a representation-based approach to mitigate it. Specifically, we first construct a prompt-only query by replacing the subject slot with a meaningless subject such as "[MASK]". By leveraging the prompt-only query, we can estimate the PLM’s bias toward the prompt, from which we generate a “biased representation”. Subsequently, we utilize the biased representation to mitigate prompt bias by vector operations in the representation layer. We conduct experiments across different prompts, PLMs and benchmarks, and find that our approach effectively rectifies the issue of inflated performance caused by prompt bias overfitting, thereby enhancing the reliability of factual knowledge benchmarks. Furthermore, our debiasing approach consistently and significantly improves the retrieval capability of prompts.
@@
  • We propose an approach to quantify the bias of knowledge-extraction prompts and assess its impact on the reliability of benchmarks. We show that all kinds of prompts in the experiments exhibit non-negligible bias, with gradient-based prompts playing significantly higher levels of bias. Furthermore, we demonstrate two negative impacts of prompt bias: 1) unreasonably amplifying benchmark performance through overfitting; 2) impairing the prompt retrieval capability by misleading PLMs.
@@
2.   Investigating Prompt Bias in Factual Knowledge Extraction
2.1.   Uncovering Prompt Bias in PLMs
@@
2.2.   Quantifying Prompt Bias in PLMs
@@
2.3.   The Impact of Prompt Bias on Factual Knowledge Extraction
@@ 4.2.   Prompt Settings
LAMA Manual refers to a series of manual prompts constructed by Petroni et al. (2019).
LPAQA refers to the mining-based and paraphrasing-based prompts constructed by Jiang et al. (2020).
@@ 4.4.   Querying Paradigms
Prec. Prec.${}^{d}$ Prec. Prec.${}^{d}$ Prec. Prec.${}^{d}$
@@ 6.   Analysis and Discussion
Datasets Prompts Prec. Prec.${}^{d}$
@@ 7.   Related Work
Bias in PLMs is widely investigated in the NLP field, from training corpus (Kurita et al., 2019; Webster et al., 2020; Dev et al., 2020) to downstream tasks such as PLM-based metrics (Sun et al., 2022), machine translation (Stanovsky et al., 2019; Prates et al., 2020; Wang et al., 2022a). Many works explore how to mitigate the intrinsic bias in PLMs (Qian et al., 2019; Bordia and Bowman, 2019; Webster et al., 2020; Qian et al., 2021; Fei et al., 2023). Counterfactual data augmentation (CDA) (Zmigrod et al., 2019; Dinan et al., 2020; Webster et al., 2020; Barikeri et al., 2021) involves re-balancing the training· corpus by swapping bias attribute words and taking further training. Karimi Mahabadi et al. (2020) and Utama et al. (2020) adjust the model’s training loss to mitigate bias by down-weighting the biased data in the corpus. Webster et al. (2020) propose using dropout regularization (Srivastava et al., 2014) as a bias mitigation approach. Schick et al. (2021) propose a post-hoc debiasing technique to discourage PLMs from generating biased text by leveraging their internal knowledge. There are also other bias mitigation technologies such as projection-based debiasing (Ravfogel et al., 2020; Liang et al., 2020), and contrastive learning debiasing (Lyu et al., 2022). In contrast to our method, most of these approaches require the extra cost of data manipulations or model retraining.
@@
Although generative language models Touvron et al. (2023); Achiam et al. (2023) have shown significant success in various language understanding and generation tasks Zhong et al. (2023a); Peng et al. (2023); Lu et al. (2023), recent studies Zheng et al. (2024); Lyu et al. (2024) show that LLMs tend to make biased choices that are inconsistent with their inherent (generated) knowledge. Our debiasing strategy consistently enhances performance and reliability on Llama2, demonstrating the universality of our method and its potential to serve as the standard post-processing toolkit for any language model. In future work, it is also worth investigating the prompt bias in other LLM-prompting-based tasks., such as the lexical choice bias in translation Ding et al. (2021), decision bias in healthcare copilot Ren et al. (2024).
@@ 8.   Conclusions
In this paper, we propose a method to quantify the prompt bias in factual knowledge extraction for various prompts and PLMs. We then demonstrate two negative impacts of prompt bias: overfitting benchmarks and misleading language models. Based on these findings, we propose an effective approach to mitigate the prompt bias using the representation vector of prompt-only querying. Experiments show that our approach effectively rectified the inflated benchmark performance achieved by prompt bias overfitting while improving the retrieval capability of prompts. Upon mitigating prompt bias, we observe close performance between state-of-the-art prompts and vanilla manual prompts. This suggests that these “better” prompts achieve their performance mainly by better overfitting test datasets via prompt bias rather than their knowledge retrieval capability. Our findings shed new light on the underlying predicting mechanisms of prompt-based queries in PLMs. Although our debiasing method is proposed for factual knowledge extraction, this approach can also be applied to other prompt-based tasks where the prompt bias occurs. For future work intending to find prompts with better retrieval capability, we strongly recommend they evaluate performance with the debiasing method.

### v2

@@ Abstract
Recent research shows that pre-trained language models (PLMs) suffer from “prompt bias” in factual knowledge extraction, i.e., prompts tend to introduce biases toward specific labels. Prompt bias presents a significant challenge in assessing the factual knowledge within PLMs. Therefore, this paper aims to improve the reliability of existing benchmarks by thoroughly investigating and mitigating prompt bias. We show that: 1) all prompts in the experiments exhibit non-negligible bias, with gradient-based prompts like AutoPrompt and OptiPrompt displaying significantly higher levels of bias; 2) prompt bias can amplify benchmark accuracy unreasonably by overfitting the test datasets, especially on imbalanced datasets like LAMA. Based on these findings, we propose a representation-based approach to mitigate the prompt bias during inference time. Specifically, we first estimate the biased representation using prompt-only querying, and then remove it from the model’s internal representations to generate the debiased representations, which are used to produce the final debiased outputs. Experiments across various prompts, PLMs, and benchmarks show that our approach can not only correct the overfitted performance caused by prompt bias, but also significantly improve the prompt retrieval capability (up to 10% absolute performance gain). These results indicate that our approach effectively alleviates prompt bias in knowledge evaluation, thereby enhancing the reliability of benchmark assessments. Hopefully, our plug-and-play approach can be a golden standard to strengthen PLMs toward reliable knowledge bases. Code and data are released in https://github.com/FelliYang/PromptBias.
@@
Ziyang Xu1, Keqin Peng3, Liang Ding4thanks:   Corresponding Authors: Liang Ding and Xiliang Lu., Dacheng Tao5, Xiliang Lu1,2,6∗
1Institute of Artificial Intelligence, School of Computer Science, Wuhan University, China
2School of Mathematics and Statistics, Wuhan University, China
6Hubei Key Laboratory of Computational Science, Wuhan University, China
3Beihang University, 4The University of Sydney, 5Nanyang Technological University
@@ 1.   Introduction
However, recent research found that the outputs of prompt-based querying are dominated by prompt bias rather than PLMs’ internal knowledge Cao et al. (2021), which strongly questions the reliability of current factual benchmarks. The interference of prompt bias makes it challenging to evaluate the amount of factual knowledge inside PLMs, significantly hindering language models from serving as reliable knowledge bases.
This paper aims to improve the reliability of factual knowledge benchmarks by thoroughly conducting a comprehensive analysis of prompt bias impact and mitigating it in the knowledge retrieval process. In this paper, we propose a general method to quantify the prompt bias across various PLMs and prompt types, and thoroughly assess its impact on widely used benchmarks such as LAMA. Experiments show that all current knowledge-extraction prompts have significant prompt bias, with previously reported high-performing prompts such as AutoPrompt and Optiprompt often exhibiting more pronounced bias. Moreover, we find that prompt bias severely compromises benchmark reliability by overfitting the test datasets and amplifying performance unreasonably. For example, the prompt bias of OptiPrompt helps inflate absolute accuracy by over 16.47% on the LAMA benchmark when probing the BERT-base model. Additionally, prompt bias can mislead language models and discourage models from making correct predictions, as illustrated in Figure 1. This significantly hinders the knowledge retrieval capabilities of prompts.
Based on these findings, we propose a representation-based approach to mitigate the prompt bias. Specifically, we first construct a prompt-only query by replacing the subject slot with a meaningless subject such as "[MASK]". By leveraging the prompt-only query, we can estimate the PLM’s bias toward the prompt, from which we generate a “biased representation”. Subsequently, we utilize the biased representation to mitigate prompt bias by vector operations in the representation layer.
We conduct experiments across different prompts, PLMs and benchmarks, and find that our approach effectively rectifies the issue of inflated performance caused by prompt bias overfitting, thereby enhancing the reliability of factual knowledge benchmarks. Furthermore, our debiasing approach consistently and significantly improves the retrieval capability of prompts.
@@
  • We propose an approach to quantify the prompt bias in knowledge extraction and assess its impact on the reliability of benchmarks. We show that all kinds of prompts in the experiments exhibit non-negligible prompt bias, with gradient-based prompts playing significantly higher levels of bias. Furthermore, we demonstrate two negative impacts of prompt bias: 1) unreasonably amplifying benchmark performance through overfitting; 2) impairing the prompt retrieval capability by misleading PLMs.
@@
2.   Investigating Prompt Bias
2.1.   Uncovering Prompt Bias
@@
2.2.   Quantifying Prompt Bias
@@
2.3.   Prompt Bias Impact on Benchmarks
@@ 4.2.   Prompt Settings
LAMA Manual refers to a series of manual prompts constructed by Petroni et al. (2019) to probe factual knowledge within PLMs. To extract factual knowledge from PLMs, Petroni et al. (2019) manually constructed a specific prompt for each relation in the LAMA benchmark. For example, the prompt “[X] was born in [Mask].” is designed for the relation place-of-birth.
LPAQA refers to the mining-based and paraphrasing-based prompts constructed by Jiang et al. (2020), which demonstrate better performance than LAMA Manual in the LAMA benchmark.
@@ 4.4.   Querying Paradigms
Prec. Prec.d Prec. Prec.d Prec. Prec.d
@@ 6.   Analysis and Discussion
Datasets Prompts Prec. Prec.d
@@ 7.   Related Work
Bias in PLMs is widely investigated in the NLP field, from training corpus (Kurita et al., 2019; Webster et al., 2020; Dev et al., 2020) to downstream tasks such as PLM-based metrics (Sun et al., 2022), machine translation (Stanovsky et al., 2019; Prates et al., 2020; Wang et al., 2022a). Many works explore how to mitigate the intrinsic bias in PLMs (Qian et al., 2019; Bordia and Bowman, 2019; Webster et al., 2020; Qian et al., 2021; Fei et al., 2023). and the extrinsic bias in downstream tasks(Zhao et al., 2017, 2018; Sun et al., 2022; Wang et al., 2022a; Behnke et al., 2022). Focusing on the intrinsic bias, counterfactual data augmentation (CDA) (Zmigrod et al., 2019; Dinan et al., 2020; Webster et al., 2020; Barikeri et al., 2021) involves re-balancing the training· corpus by swapping bias attribute words and taking further training. Karimi Mahabadi et al. (2020) and Utama et al. (2020) adjust the model’s training loss to mitigate bias by down-weighting the biased data in the corpus. Webster et al. (2020) propose using dropout regularization (Srivastava et al., 2014) as a bias mitigation approach. Schick et al. (2021) propose a post-hoc debiasing technique to discourage PLMs from generating biased text by leveraging their internal knowledge. There are also other bias mitigation technologies such as projection-based debiasing (Ravfogel et al., 2020; Liang et al., 2020), and contrastive learning debiasing (Lyu et al., 2022). In contrast to our method, most of these approaches require the extra cost of data manipulations or model retraining.
@@
Although generative language models Touvron et al. (2023); Achiam et al. (2023) have shown significant success in various language understanding and generation tasks Zhong et al. (2023a); Peng et al. (2023); Lu et al. (2023), recent studies Zheng et al. (2024); Lyu et al. (2024) show that LLMs tend to make biased choices that are inconsistent with their inherent (generated) knowledge. Our debiasing strategy consistently enhances performance and reliability on Llama2, demonstrating the universality of our method and its potential to serve as the standard post-processing toolkit for large language model. In future work, it is also worth investigating the prompt bias in other LLM-prompting-based tasks, such as the lexical choice bias in translation Ding et al. (2021), decision bias in healthcare copilot Ren et al. (2024).
@@ 8.   Conclusions
In this paper, we propose a method to quantify the prompt bias in factual knowledge extraction and demonstrate two negative impacts of prompt bias: overfitting benchmarks and misleading language models. Based on these findings, we propose an approach to mitigate the prompt bias using the representation vector of prompt-only querying. Experiments show that our approach effectively rectified the inflated benchmark performance achieved by prompt bias overfitting, resulting in a more reliable evaluation of factual knowledge within PLMs. Furthermore, our approach significantly enhances the retrieval performance of prompts, even within contemporary large language models like Llama2. Additionally, we observe close performance between state-of-the-art prompts and vanilla manual prompts after mitigating prompt bias. This suggests that these “better” prompts achieve their performance mainly by better overfitting test datasets via prompt bias rather than their knowledge retrieval capability. These findings contribute to the development of PLM-based knowledge base construction by shedding new light on the reliability of existing benchmarks and the actual amount of factual knowledge within PLMs. Although our debiasing method is proposed for factual knowledge extraction, this approach can also be applied to other prompt-based tasks where the prompt bias occurs such as topic classification. For future work intending to find prompts with better retrieval capability, we strongly recommend they evaluate performance with the debiasing method.